# Convergence of spectral discretization for the flow of diffeomorphisms

### v1

@@ 2 The Riemannian manifold of Sobolev diffeomorphisms
    for an inner product $(\!(v,w)\!)_{H^{m}}=\langle\mathcal{L}v,w\rangle$ with $\mathcal{L}=B^{*}B+\bar{\mathcal{L}}$ for $B$ a differential operator of order $m$ with bounded coefficients and compact self-adjoint ¯$\bar{\mathcal{L}}:H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\to H^{-m}(\mathbb{T}^{d};\mathbb{R}^{d})$, continuous into $H^{-m+1}(\mathbb{T}^{d};\mathbb{R}^{d})$, geodesics satisfy a weak PDE, whose strong form is (6) [GRRW23].
@@ 3 Refined and new regularity estimates
In this section we give a regularity result for geodesics, theorem 11, that allows to prove convergence of the spectral space discretization. Furthermore, in order to be able to provide the explicit dependence of the final error estimates on the initial condition $v_{0}$ of the geodesic, we revisit and refine in lemmas 7 and 8 a few known regularity estimates for operations with Sobolev diffeomorphisms. For completeness we also provide in lemma 10 a sufficient condition for $l$-fold differentiability of the Riemannian metric on $\mathcal{D}_{\mathrm{id}}^{m}$, since our convergence results require two- and threefold differentiability. Again the results and proofs can readily be adapted from the torus $\mathbb{T}^{d}$ to other domains such as $\mathbb{R}^{d}$ or smooth compact manifolds.
We begin with an estimate for the composition with a small deformation (iterations of which will then yield estimates for general diffeomorphisms), which refines the argument in [IKT13, Lemma 2.7]. Below we will abbreviate the closed norm ball in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ of radius $\varepsilon>0$ by
@@
for some $\hat{K},K>0$ by [IKT13, Lemma 2.3] (essentially since $H^{m-1}(\mathbb{T}^{d})$ is a Banach algebra). In summary,
@@
From now on, $v$ is an arbitrary velocity field. We split it into $N$ time segments $v^{j}=v|_{[t_{j-1},t_{j}]}$ (extended to $I$ by zero) with $10=t_{0}<\ldots<t_{N}=$ 1 chosen such that $\int_{t_{j-1}}^{t_{j}}\|v_{t}\|_{H^{m}}\mathrm{d}t=\|v\|_{L^{1}(I;H^{m})}/N<\frac{\varepsilon}{1+C\varepsilon}$. Let us denote by $\phi^{j}$ the flow of $v^{j}$, then by the above we have $\|\phi^{j}_{t}-\mathrm{id}\|_{H^{m}}\leq\varepsilon$, and furthermore
@@
    using [BH21, Thm. 6.1] with constant $c>0$ in the last inequality and lemma 7 in the first. The latter is allowed since $(\phi_{t}^{j})^{-1}$ is the flow of $[0,t]\ni r\mapsto-v^{j}(t-r)$ [Tro95] so that by (7) we also have $\|(\phi_{t}^{j})^{-1}-\mathrm{id}\|_{H^{m}}\leq V_{N}<\varepsilon$. Consequently, with some $\tilde{c}>0$ we have $\|\mathrm{Ad}_{\phi_{t}^{j}}w\|_{H^{s}}\leq(1+CV_{N})(1+\tilde{c}V_{N})\|w\|_{H^{s}}$, as desired. ∎
We will also need certain continuity properties of composition and adjoint map (note that continuity of $H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})\times\mathcal{D}_{\mathrm{id}}^{m}\ni(w,\phi)\mapsto(w\circ\phi,\mathrm{Ad}_{\phi}w)\in H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})\times H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})$ for $m>\frac{d}{2}+1$ and $0\leq s<m$ is known [IKT13, Lemma 2.7]).
@@
since $\eta\circ\phi_{n}^{-1}\to\eta\circ\phi^{-1}$ strongly in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ (by continuity of inversion in $\mathcal{D}_{\mathrm{id}}^{m}$ and composition with $\mathcal{D}_{\mathrm{id}}^{m}$ [IKT13, Lemmas 2.7-2.8]), $\det D\phi_{n}^{-1}\to\det D\phi^{-1}$ strongly in $H^{m-1}(\mathbb{T}^{d})$ (since $\phi\mapsto\phi^{-1}$ is continuous on $\mathcal{D}_{\mathrm{id}}^{m}$, $\psi\mapsto D\psi$ is continuous from $\mathcal{D}_{\mathrm{id}}^{m}$ to $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d\times d})$, and taking the pointwise determinant is continuous from $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d\times d})$ into $H^{m-1}(\mathbb{T}^{d})$ since $H^{m-1}$ forms a Banach algebra [IKT13, Lemma 2.16]), and thus $\det D\phi_{n}^{-1}\,\eta\circ\phi_{n}^{-1}\to\det D\phi^{-1}\,\eta\circ\phi^{-1}$ strongly in $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d})$ by [BH21, Thm. 6.1].
@@
Before proving a regularity estimate for geodesics (which then will allow an approximation via spectral discretization), we need to ensure well-posedness of geodesics in the first place. Even though shortest geodesics exist between any two points of $\mathcal{D}_{\mathrm{id}}^{m}$ [Tro95, BV17] and under certain conditions on the inner product $(\!(\cdot,\cdot)\!)_{H^{m}}$ even satisfy a (weak) geodesic equation [GRRW23], the solution of this geodesic equation forward in time may not be unique. To this end we appeal to classical differential geometry: If the Riemannian metric $g_{\phi}(v,w)=\langle\mathcal{L}(v\circ\phi),w\circ\phi\rangle$ is twice continuously differentiable in $\phi$, then the right-hand side of the second order geodesic equation is Lipschitz (even differentiable) and thus locally uniquely solvable [Kli95, § 1.6] with the solution depending continuously on the initial condition by classical ODE theory. In other words, the Riemannian exponential map is locally well-defined and continuous (and $n$ times differentiable if the metric has $n$ further derivatives). For the sake of completeness we provide a sufficient condition for this differentiability requirement, which essentially is a variation of [MeP10, Thm. 4.1].
@@
Let $m>\frac{d}{2}+1$, $0\leq l\leq m$. If =¯$\mathcal{L}=\bar{\mathcal{L}}+B^{*}B$ with ¯$\bar{\mathcal{L}}:H^{m-l}(\mathbb{T}^{d};\mathbb{R}^{d})\to H^{l-m}(\mathbb{T}^{d};\mathbb{R}^{d})$ bounded linear and $B=\sum_{k=0}^{m}a_{k}D^{k}$ a differential operator of order $m$ with coefficient tensors $a_{i}$ of $W^{l,\infty}$-regularity, then the metric (3) is $l$ times continuously differentiable.
@@
Now $(w,\phi)\mapsto w\circ\phi^{-1}$ is $l$ times differentiable as a map from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ to $H^{m-l}(\mathbb{T}^{d};\mathbb{R}^{d})$. Indeed, it is linear in $w$ so that the map and all its derivatives with respect to $\phi$ are $C^{\infty}$-smooth in $w$ as long as they are continuous in $w$. Thus it remains to check differentiability with respect to $\phi$. From the formulas
@@
we see inductively that the $l$th derivative of $v\circ\phi^{-1}$ with respect to $\phi$ in directions $u_{1},\ldots,u_{l}\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ is a sum of products of $D\phi^{-1}\in H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d\times d})$ with derivatives of $\phi^{-1}$, $v$, and $u_{1},\ldots,u_{l}$ (all derivatives composed with $\phi^{-1}$ except for the derivatives of $\phi^{-1}$). Moreover, these derivatives amount to an overall order of $l+1$ in each product, with the highest occurring order not exceeding $l$. By [BH21, Thm. 6.1], the product of functions in $H^{m-n_{0}},\ldots,H^{m-n_{l}}$ with $n_{0},\ldots,n_{k}\geq 1$ and $n_{k+1}=\ldots=n_{l}=0$ is bounded in $H^{m+k(m-\frac{d}{2})-n_{0}-\ldots-n_{k}}$; in particular, if $n_{0}+\ldots+n_{k}=l+1$ and $k>0$ the product is bounded in $H^{m-l}$. Thus, the $l$th derivative of $v\circ\phi^{-1}$ in $\phi$ is a bounded $l$-linear map from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})^{l}$ into $H^{m-l}(\mathbb{T}^{d};\mathbb{R}^{d})$, which depends continuously on $\phi\in\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$.
@@
$\bar{D}w=[Dw\circ\phi^{-1}\,D(\phi^{-1})]\circ\phi=Dw(D\phi)^{-1},$
the map $(w,\phi)\mapsto\bar{D}w$ is infinitely smooth from $H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into $H^{s-1}(\mathbb{T}^{d})^{d\times d}$ (even though $\phi\mapsto\phi^{-1}$ is not smooth in $\mathcal{D}_{\mathrm{id}}^{m}$, since matrix inversion and pointwise multiplication is). By iteration, $(w,\phi)\mapsto\bar{D}^{m}w$ is smooth from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into $L^{2}(\mathbb{T}^{d})^{d\times\cdots\times d}$. Finally, $\phi\mapsto a_{i}\circ\phi$ is $l$-times differentiable from $\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into tensor fields of $L^{\infty}$-regularity so that $(u,\phi)\mapsto B(u\circ\phi^{-1})\circ\phi$ is differentiable from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into $L^{2}(\mathbb{T}^{d})$-tensors. Since pointwise product and integration are smooth as long as they are bounded, $(u,v,\phi)\mapsto g_{\phi}(u,v)$ is $l$ times differentiable. ∎
@@
which is still well-defined due to $v_{t}\in H^{m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$. However, this is exactly the weak form of the geodesic ODE (6) which by [GRRW23] rigorously characterizes the geodesics. Since the geodesic ODE is uniquely solvable by the regularity of the metric, $\rho_{t}$ must be the corresponding momentum. The continuous time dependence finally follows from lemma 9 and the continuity of the geodesic in $\mathcal{D}_{\mathrm{id}}^{m}$. ∎
@@
We formulated our result for $k\leq m$ since in that range the momentum $\rho_{t}$ stays a distribution. However, the argument extends to $m<k\leq m+\frac{d}{2}$. Indeed, by theorem 11 we already know $\rho_{t}=\mathrm{Ad}_{\phi_{t}^{-1}}^{*}\rho_{0}$, and it remains to show that $\rho_{0}\in H^{-m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$ implies $\rho_{t}\in H^{-m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$ for all $t$. Now let $w:\mathbb{T}^{d}\to\mathbb{R}^{d}$ be smooth, then
@@
where the last line follows from [BH21, Thm.6.1], using $m-1>\frac{d}{2}\geq k-m$. Since the determinant is a degree $d$ polynomial and $H^{m-1}$ is a Banach algebra (cf. [BH21, Thm.6.1]), we have $\|\det D\phi^{-1}\|_{H^{m-1}}\lesssim\|D\phi^{-1}\|_{H^{m-1}}^{d}\lesssim(1+\|\phi^{-1}-\mathrm{id}\|_{H^{m}})^{d}$. Furthermore, $\|\rho_{0}\circ\phi_{t}^{-1}\|_{H^{k-m}}\lesssim c(\phi_{t}^{-1})\|\rho_{0}\|_{H^{k-m}}$ by (2) so that overall

### v2

@@ 2 The Riemannian manifold of Sobolev diffeomorphisms
    for an inner product $(\!(v,w)\!)_{H^{m}}=\langle\mathcal{L}v,w\rangle$ with $\mathcal{L}=B^{*}B+\bar{\mathcal{L}}$ for $B$ a differential operator of order $m$ with bounded coefficients and compact self-adjoint $\bar{\mathcal{L}}:H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\to H^{-m}(\mathbb{T}^{d};\mathbb{R}^{d})$, continuous into $H^{-m+1}(\mathbb{T}^{d};\mathbb{R}^{d})$, geodesics satisfy a weak PDE, whose strong form is (6) [GRR+23] (cf. lemma 14).
@@ 3 Refined and new regularity estimates
In this section we give a regularity result for geodesics, theorem 15, that allows to prove convergence of the spectral space discretization. Furthermore, in order to be able to provide the explicit dependence of the final error estimates on the initial condition $v_{0}$ of the geodesic, we revisit and refine in lemmas 9 and 10 a few known regularity estimates for operations with Sobolev diffeomorphisms. For completeness we also provide in lemma 12 a sufficient condition for $l$-fold differentiability of the Riemannian metric on $\mathcal{D}_{\mathrm{id}}^{m}$, since our convergence results require two- and threefold differentiability. Again the results and proofs can readily be adapted from the torus $\mathbb{T}^{d}$ to other domains such as $\mathbb{R}^{d}$ or smooth compact manifolds.
We first recall a continuity result from [BH21] (there given for general Lipschitz domains and integrability exponents) for pointwise multiplication of Sobolev functions that we will frequently make use of.
Lemma 7 (Continuity of multiplication, [BH21, Cor. 6.3, Thm. 7.4]).
Let $0\leq s\leq r,m$ with $r+m>s+\frac{d}{2}$, then pointwise multiplication is a continuous bilinear map $H^{m}(\mathbb{T}^{d})\times H^{r}(\mathbb{T}^{d})\to H^{s}(\mathbb{T}^{d})$.
We begin with an estimate for the composition with a small deformation (iterations of which will then yield estimates for general diffeomorphisms), which refines the following statement from [IKT13] (given there on $\mathbb{R}^{d}$ instead of $\mathbb{T}^{d}$, but the proof carries over).
Lemma 8 (Continuity of composition, [IKT13, Lemma 2.7]).
Let $m>\frac{d}{2}+1$ and $0\leq s\leq m$. Then $H^{s}(\mathbb{T}^{d})\times\mathcal{D}^{m}\ni(f,\phi)\mapsto f\circ\phi\in H^{s}(\mathbb{T}^{d})$ is continuous. Moreover, for any $m,M>0$ there exists $C>0$ such that $\det D\phi\geq m$ and $\|\phi-\mathrm{id}\|_{H^{m}}\leq M$ for $\phi\in\mathcal{D}^{m}$ imply
$\|f\circ\phi\|_{H^{s}}\leq C\|f\|_{H^{s}}\qquad\text{for all }f\in H^{s}(\mathbb{T}^{d}).$
Our refinement consists in the precise control of the involved constant in terms of how close the diffeomorphism is to the identity. Below we will abbreviate the closed norm ball in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ of radius $\varepsilon>0$ by
@@
for some $\hat{K},K>0$ by [IKT13, Lemma 2.3] or lemma 7 (essentially since $H^{m-1}(\mathbb{T}^{d})$ is a Banach algebra). In summary,
@@
From now on, $v$ is an arbitrary velocity field. We split it into $N$ time segments $v^{j}=v|_{[t_{j-1},t_{j}]}$ (extended to $I$ by zero) with $0=t_{0}<\ldots<t_{N}=1$ chosen such that $\int_{t_{j-1}}^{t_{j}}\|v_{t}\|_{H^{m}}\mathrm{d}t=\|v\|_{L^{1}(I;H^{m})}/N<\frac{\varepsilon}{1+C\varepsilon}$. Let us denote by $\phi^{j}$ the flow of $v^{j}$, then by the above we have $\|\phi^{j}_{t}-\mathrm{id}\|_{H^{m}}\leq\varepsilon$, and furthermore
@@
    using lemma 7 with constant $c>0$ in the last inequality and lemma 9 in the first. The latter is allowed since $(\phi_{t}^{j})^{-1}$ is the flow of $0,t]\ni r\mapsto-v^{j}(t-r) [$TRO95] so that by (7) we also have $\|(\phi_{t}^{j})^{-1}-\mathrm{id}\|_{H^{m}}\leq V_{N}<\varepsilon$. Consequently, with some $\tilde{c}>0$ we have $\|\mathrm{Ad}_{\phi_{t}^{j}}w\|_{H^{s}}\leq(1+CV_{N})(1+\tilde{c}V_{N})\|w\|_{H^{s}}$, as desired. ∎
We will also need certain continuity properties of composition and adjoint map (note that continuity of $H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})\times\mathcal{D}_{\mathrm{id}}^{m}\ni(w,\phi)\mapsto(w\circ\phi,\mathrm{Ad}_{\phi}w)\in H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})\times H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})$ for $m>\frac{d}{2}+1$ and $0\leq s<m$ is known, see lemma 8).
@@
since $\eta\circ\phi_{n}^{-1}\to\eta\circ\phi^{-1}$ strongly in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ (by continuity of inversion in $\mathcal{D}_{\mathrm{id}}^{m}$ and composition with $\mathcal{D}_{\mathrm{id}}^{m}$ [IKT13, Lemmas 2.7-2.8]), $\det D\phi_{n}^{-1}\to\det D\phi^{-1}$ strongly in $H^{m-1}(\mathbb{T}^{d})$ (since $\phi\mapsto\phi^{-1}$ is continuous on $\mathcal{D}_{\mathrm{id}}^{m}$, $\psi\mapsto D\psi$ is continuous from $\mathcal{D}_{\mathrm{id}}^{m}$ to $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d\times d})$, and taking the pointwise determinant is continuous from $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d\times d})$ into $H^{m-1}(\mathbb{T}^{d})$ since $H^{m-1}$ forms a Banach algebra [IKT13, Lemma 2.16]), and thus $\det D\phi_{n}^{-1}\,\eta\circ\phi_{n}^{-1}\to\det D\phi^{-1}\,\eta\circ\phi^{-1}$ strongly in $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d})$ by lemma 7.
@@
Before proving a regularity estimate for geodesics (which then will allow an approximation via spectral discretization), we need to ensure well-posedness of geodesics in the first place. Even though shortest geodesics exist between any two points of $\mathcal{D}_{\mathrm{id}}^{m}$ [TRO95, Thm. 3.7],[BV17, Thm. 7.2] [^3] Note that [BV17, Thm. 7.2] only states the result for smooth metrics, but does not use the smoothness in proving existence of path energy minimizers. and under certain conditions on the inner product $(\!(\cdot,\cdot)\!)_{H^{m}}$ even satisfy a (weak) geodesic equation [GRR+23] (cf. lemma 14), the solution of this geodesic equation forward in time may not be unique. To this end we appeal to classical differential geometry: If the Riemannian metric $g_{\phi}(v,w)=\langle\mathcal{L}(v\circ\phi),w\circ\phi\rangle$ is twice continuously differentiable in $\phi$, then the right-hand side of the second order geodesic equation is Lipschitz (even differentiable) and thus locally uniquely solvable [KLI95, § 1.6] with the solution depending continuously on the initial condition by classical ODE theory. In other words, the Riemannian exponential map is locally well-defined and continuous (and $n$ times differentiable if the metric has $n$ further derivatives). For the sake of completeness we provide a sufficient condition for this differentiability requirement, which essentially is a variation of [MP10, Thm. 4.1].
@@
Let $m>\frac{d}{2}+1$, $0\leq l\leq m$. If $\mathcal{L}=\bar{\mathcal{L}}+B^{*}B$ with $\bar{\mathcal{L}}:H^{m-l}(\mathbb{T}^{d};\mathbb{R}^{d})\to H^{l-m}(\mathbb{T}^{d};\mathbb{R}^{d})$ bounded linear and $B=\sum_{k=0}^{m}a_{k}D^{k}$ a differential operator of order $m$ with coefficient tensors $a_{i}$ of $W^{l,\infty}$-regularity, then the metric (3) is $l$ times continuously differentiable.
Remark 13 (Example metrics).
The result of course applies to the standard example $\mathcal{L}=1+(-\Delta)^{m}$ with the choice $\bar{\mathcal{L}}=0$ and $B=(1,\Delta^{m/2})$ (for even $m$) or $(1,\nabla\Delta^{(m-1)/2})$ (for odd $m$), but one may just as well introduce non-local operators or spatially inhomogeneous weights, e.g. by choosing $\bar{\mathcal{L}}u(x)=\int_{\mathbb{T}^{d}}\frac{u(x)-u(y)}{\mathrm{dist}(x,y)^{d+2s}}\omega(x,y)\,\mathrm{d}y$ for some $s\in(0,1)\cap(0,m-l)$ and a bounded weight $\omega(x,y)$ or by choosing $\bar{\mathcal{L}}u=-\mathrm{div}(\chi_{A}\nabla u)$ for the characteristic function $\chi_{A}$ of some measurable $A\subset\mathbb{T}^{d}$ (as long as $l\leq m-1$). In particular the latter choice may be of interest in applications, modelling an enhanced energy dissipation (which is one possible physical interpretation of the path energy) in region $A$, e.g. due to a low permeability porous structure. The resulting metric is not smooth, but only m1m-1 times continuously differentiable (as one can for instance easily verify for the concrete choice $d=1,m=2,A=[0,\frac{1}{2}]$).
The conditions of lemma 12 are clearly only sufficient: For instance, adding a Fourier multiplier of order $m$ or even an operator from $\mathrm{OPS}_{1,0}^{m}$ to $B$ does not change the differentiability properties of the metric, since these induce smooth metrics by [BEK15, BBC+20].
@@
Now $(w,\phi)\mapsto w\circ\phi^{-1}$ is $l$ times differentiable as a map from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ to $H^{m-l}(\mathbb{T}^{d};\mathbb{R}^{d})$. Indeed, it is linear in $w$ so that the map and all its derivatives with respect to $\phi$ are $C^{\infty}$-smooth in $w$ as long as they are continuous in $w$. Thus it remains to check differentiability with respect to $\phi$.
Using the formulas
@@
we first show inductively that $\partial_{\phi}^{l}(v\circ\phi^{-1})(u_{1},\ldots,u_{l})$ is a sum of products of (potentially multiple factors) $D\phi^{-1}$ with (potentially zero or higher order) derivatives of $v$ and $u_{1},\ldots,u_{l}$ (all composed with $\phi^{-1}$) and (potentially multiple, first or higher order) derivatives of $D\phi^{-1}$. Moreover, these derivatives amount to an overall order of $l$ in each product, at most l1l-1 of which can act on $D\phi^{-1}$. Indeed, this holds for $l=1$, and if it holds for $l$, then applying another time $\partial_{\phi}$ in some direction $u_{l+1}$ leads by the product rule again to a sum of terms, each one resulting from differentiating one factor of the previous products with respect to $\phi$. If this factor is $D^{j}w\circ\phi^{-1}$ for $w$ being one of $v,u_{1},\ldots,u_{l}$, then
$\partial_{\phi}[D^{j}w\circ\phi^{-1}](u_{l+1})=-D^{j+1}w\circ\phi^{-1}\,D\phi^{-1}\,u_{l+1}\circ\phi^{-1}$
so that the total order of derivatives of $v,u_{1},\ldots,u_{l+1},D\phi^{-1}$ indeed was increased by one (from $l$ to $l+1$). If the factor is $D^{j}(D\phi^{-1})$, then
$\partial_{\phi}[D^{j}(D\phi^{-1})](u_{l+1})=-D^{j+1}(D\phi^{-1}\,u_{l+1}\circ\phi^{-1}),$
which by the product rule and the same arguments as above increases the total order of derivatives of $v,u_{1},\ldots,u_{l+1},D\phi^{-1}$ again by one (and the highest occurring order of differentiating $D\phi^{-1}$ at most by one), as desired.
In summary, for $l\geq 1$ we have that $\partial_{\phi}^{l}(v\circ\phi^{-1})(u_{1},\ldots,u_{l})$ is a sum of products with one factor in each of $H^{m-n_{1}},\ldots,H^{m-n_{k}}$ (the differentiated $v$ and $u_{j}$), one factor in each of $H^{m-1-n_{k+1}},\ldots,H^{m-1-n_{K}}$ (the differentiated $D\phi^{-1}$) as well as (potentially multiple) factors in $H^{m}$ (undifferentiated $u_{j}$) and $H^{m-1}$ (the factors $D\phi^{-1}$), where $n_{1},\ldots,n_{K}\geq 1$ with $n_{1}+\ldots+n_{K}=l$ and $n_{k+1},\ldots,n_{K}\leq l-1$. Now lemma 7 implies that as soon as one function has Sobolev regularity larger than $\frac{d}{2}$, the product of two functions has the minimum Sobolev regularity of the two. So assume first $m-n_{1},\ldots,m-n_{k},m-1-n_{k+1},\ldots,m-1-n_{K}>\frac{d}{2}$, then the product has Sobolev regularity $\min\{m-n_{1},\ldots,m-n_{k},m-1-n_{k+1},\ldots,m-1-n_{K}\}\geq m-l$. Otherwise, if at least one factor has Sobolev regularity no larger than $\frac{d}{2}$, all factors with Sobolev regularity larger than $\frac{d}{2}$ can be ignored for the regularity of the product. So assume without loss of generality that $m-n_{1},\ldots,m-n_{k},m-1-n_{k+1},\ldots,m-1-n_{K}\leq\frac{d}{2}$ (if this only holds for a subset, then reindex and redefine $k,K$ correspondingly). Furthermore, we may assume $k=1$ without loss of generality (since for fixed $n_{1},\ldots,n_{K}$ this represents the case of overall lowest Sobolev regularity). Then, by iteratively applying lemma 7 (starting with the most regular function and multiplying one after another the functions in the order of decreasing Sobolev regularity) we obtain that the product has Sobolev regularity $m-n_{1}-\ldots-n_{K}+(K-1)(m-1-\frac{d}{2})-\epsilon$ for any $\epsilon>0$ (note that lemma 7 also holds for fractional exponents). Due to $m>1+\frac{d}{2}$ and $n_{1}+\ldots+n_{K}\leq l$ this implies that the product has Sobolev regularity mlm-l. Hence in total, $\partial_{\phi}^{l}(v\circ\phi^{-1})(u_{1},\ldots,u_{l})\in H^{m-l}$, where the dependence on $\phi,v,u_{1},\ldots,u_{l}\in H^{m}$ is continuous; consequently, the $l$th derivative of $v\circ\phi^{-1}$ in $\phi$ is a bounded $l$-linear map from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})^{l}$ into $H^{m-l}(\mathbb{T}^{d};\mathbb{R}^{d})$, which depends continuously on $\phi\in\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$.
@@
$\bar{D}w=[Dw\circ\phi^{-1}\,D\phi^{-1}]\circ\phi=Dw(D\phi)^{-1},$
the map $(w,\phi)\mapsto\bar{D}w$ is infinitely smooth from $H^{s}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into $H^{s-1}(\mathbb{T}^{d})^{d\times d}$ (even though $\phi\mapsto\phi^{-1}$ is not smooth in $\mathcal{D}_{\mathrm{id}}^{m}$, since matrix inversion and pointwise multiplication is). By iteration, $(w,\phi)\mapsto\bar{D}^{m}w$ is smooth from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into $L^{2}(\mathbb{T}^{d})^{d\times\cdots\times d}$. Finally, $\phi\mapsto a_{i}\circ\phi$ is $l$-times differentiable from $\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into tensor fields of $L^{\infty}$-regularity so that $(u,\phi)\mapsto B(u\circ\phi^{-1})\circ\phi$ is $l$ times differentiable from $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\times\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}$ into $L^{2}(\mathbb{T}^{d})$-tensors. Since pointwise product and integration are smooth as long as they are bounded, $(u,v,\phi)\mapsto g_{\phi}(u,v)$ is $l$ times differentiable. ∎
@@
For reference, we also recall the result from [GRR+23] on the weak PDE solved by geodesics (there formulated for a domain $D\subset\mathbb{R}^{d}$), which is based on a similar decomposition of $\mathcal{L}$.
Lemma 14 (Weak geodesic PDE, [GRR+23, Thm. 1]).
Let $m>1+\frac{d}{2}$ and the Riemannian metric (3) on $\mathcal{D}^{m}$ be induced by $\mathcal{L}=B^{*}B+\bar{\mathcal{L}}$ with $\bar{\mathcal{L}}:H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})\to H^{1-m}(\mathbb{T}^{d};\mathbb{R}^{d})$ bounded linear and $B=\sum_{k=0}^{m}a_{k}D^{k}$ a differential operator of order $m$ with coefficient tensors $a_{i}$ of $W^{0,\infty}$-regularity. Then geodesics in $\mathcal{D}^{m}$ with velocity $v$ and momentum $\rho$ satisfy the weak PDE
$0=\int_{0}^{1}\!\!-2\langle\bar{\mathcal{L}}v_{t},Dv_{t}\eta_{t}\rangle+2\langle\rho_{t},\dot{\eta}_{t}+D\eta_{t}v_{t}\rangle+\int_{\mathbb{T}^{d}}|Bv_{t}|^{2}\mathrm{div}\eta_{t}+2Bv_{t}\cdot{\mathcal{U}}(v_{t},\eta_{t})\,\mathrm{d}x\,\mathrm{d}t$
for all smooth $\eta:[0,1]\times\mathbb{T}^{d}\to\mathbb{R}^{d}$, where
${\mathcal{U}}(v_{t},\eta_{t})=\frac{\mathrm{d}}{\mathrm{d}\epsilon}\left[B(v_{t}\circ(\mathrm{id}+\epsilon\eta_{t})^{-1})\circ(\mathrm{id}+\epsilon\eta_{t})\right]_{\epsilon=0}\in H^{0}((0,1)\times\mathbb{T}^{d};\mathbb{R}^{d}).$
Its strong form is (6).
In fact, the proof does not require the specific form of $B$, but only uses differentiability of $H^{m}\times\mathcal{D}^{m}\ni(v,\phi)\mapsto[B(v\circ\phi^{-1})]\circ\phi\in L^{2}$.
@@
Note that for an $m+2$ times differentiable Riemannian metric (and thus $m$ times differentiable Riemannian exponential map) this regularity preservation result already follows from [BRU17, main theorem and Sec. 5.2]. In contrast to our direct computation, there the proof is based on an abstract argument trading in differentiability of the Riemannian exponential for preservation of Sobolev exponents, exploiting the right-invariance of the metric.
@@
which is still well-defined due to $v_{t}\in H^{m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$. However, this is exactly the weak form of the geodesic ODE (6) which by lemma 14 rigorously characterizes the geodesics. Since the geodesic ODE is uniquely solvable by the regularity of the metric, $\rho_{t}$ must be the corresponding momentum. The continuous time dependence finally follows from lemma 11 and the continuity of the geodesic in $\mathcal{D}_{\mathrm{id}}^{m}$. ∎
@@
We formulated our result for $k\leq m$ since in that range the momentum $\rho_{t}$ stays a distribution. However, the argument extends to $m<k\leq m+\frac{d}{2}$ (and the result is known to hold for $k>m+\frac{d}{2}$, see later). Indeed, by theorem 15 we already know $\rho_{t}=\mathrm{Ad}_{\phi_{t}^{-1}}^{*}\rho_{0}$, and it remains to show that $\rho_{0}\in H^{-m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$ implies $\rho_{t}\in H^{-m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$ for all $t$. Now let $w:\mathbb{T}^{d}\to\mathbb{R}^{d}$ be smooth, then
@@
where the last line follows from lemma 7, using $m-1>\frac{d}{2}\geq k-m$. Since the determinant is a degree $d$ polynomial and $H^{m-1}$ is a Banach algebra (cf. lemma 7), we have $\|\det D\phi^{-1}\|_{H^{m-1}}\lesssim\|D\phi^{-1}\|_{H^{m-1}}^{d}\lesssim(1+\|\phi^{-1}-\mathrm{id}\|_{H^{m}})^{d}$. Furthermore, $\|\rho_{0}\circ\phi_{t}^{-1}\|_{H^{k-m}}\lesssim c(\phi_{t}^{-1})\|\rho_{0}\|_{H^{k-m}}$ by (2) so that overall